OpenStudy (anonymous):

I WILL MEDAL/FAN Please help me! (NO ANSWERS PLEASE) :) Bella plans to put $200 into a savings account. She can place her money into an account represented by p(x) = 3x + 200, or into another account represented by n(x) = 200(1.04)x. Which account has the highest value in 4 years? Which account has the highest in 15 years?

OpenStudy (anonymous):

p(x) has the highest value in 4 years; p(x) has the highest value in 15 years n(x) has the highest value in 4 years; p(x) has the highest value in 15 years n(x) has the highest value in 4 years; n(x) has the highest value in 15 years p(x) has the highest value in 4 years; n(x) has the highest value in 15 years

OpenStudy (anonymous):

@doc.brown

OpenStudy (anonymous):

@CaptainLlama49

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

An easier way to get someone to look at it for you is that when you post the question, type in big letters: "MEDAL WILL BE AWARDED" or "WILL MEDAL" or something like that. I know it works with me almost always. Why don't you try reposting the question?

OpenStudy (anonymous):

Trust me, It Really Helps

OpenStudy (anonymous):

@Clueless_much78

OpenStudy (anonymous):

\[p(x) = 3x + 200\\p(4)=3\times 4+200\] \[n(x) = 200(1.04)^x\\ n(4)=200\times (1.04)^4\] compare and contrast as the teachers say

OpenStudy (anonymous):

wait what? @satellite73

OpenStudy (anonymous):

Try asking your question on the Geometry subject instead

OpenStudy (anonymous):

plug in the numbers, see what you get

OpenStudy (anonymous):

Or that...

OpenStudy (anonymous):

@Tide_Poole can you help me understand (I failed algebra 1 twice, I am taking it over, and stil confused)

OpenStudy (anonymous):

It's been quite awhile since I've done Algebra omg

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

lol its a-ok...:) @Tide_Poole

OpenStudy (anonymous):

@satellite73 `

OpenStudy (anonymous):

He didn't understand what you told him last time

OpenStudy (anonymous):

I did not :( xD I am so slow to catch on, you guys might as well try to explain this to a blind turtle.

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

I totally understand the feeling

OpenStudy (anonymous):

you have two functions, right? one is \[p(x) = 3x + 200,\]

OpenStudy (anonymous):

the other is \[ n(x) = 200(1.04)^x\]

OpenStudy (anonymous):

you are asked "Which account has the highest value in 4 years? "

OpenStudy (anonymous):

to find the answer, where you see an \(x\) replace it by a \(4\) the first one to compute is \[p(4)=3\times 4+200\] what is that number?

OpenStudy (anonymous):

did i lose you there?

OpenStudy (anonymous):

I'm not sure. He pops in and out all the time

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Lets try calling him

OpenStudy (anonymous):

@Clueless_much78

OpenStudy (anonymous):

p4=212 sorry My mum called me downstairs for a min... :( @Tide_Poole @satellite73

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

That's understandable

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

Stupid summer, xD no one can help meh :(

OpenStudy (anonymous):

@hartnn can you help meh please? (I do NOT want an answer!) :)

OpenStudy (anonymous):

@ganeshie8 @dan815 @Australopithecus

OpenStudy (anonymous):

sorry if that was counted as mass- tagging :(

OpenStudy (australopithecus):

Look at what satellite73 said. Just plug the numbers into both functions and whatever gives you the largest number is the bank account to go with

OpenStudy (anonymous):

I did that and concluded that C would be the answer? @Australopithecus

OpenStudy (australopithecus):

that is correct.